OpenStudy (anonymous):

2x-8y divided by x^2-16y^2 factor the numerator and denominator and simplify

OpenStudy (anonymous):

numerator=2(x-4y) denominator=(x+4y)(x-4y) on simplifying u get, 2/(x+4)

OpenStudy (anonymous):

oops! missed y , 2/(x+4y)

OpenStudy (anonymous):

2(x-4y)/(x+4y)(x-4y),, the (x-4y) ancel out, to simplify 2/(x+4y)